# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ IV-B Tangent Markers in An Arc
Proposition 3 and Proposition 4 analyzed the feasibility of an arc near the end points. An arc might be tangent to some constrained boundary, where light perturbations may affect the feasibility of the arc. The following proposition provides a sufficient and necessary condition for the feasibility of the optimal trajectory tangent to the constrained boundary.
@@
Assume that ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is an arc for $t\in\left[t_{0},t_{2}\right]$. ${\mathcal{S}}$ is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq 0$ at $t_{1}\in\left(t_{0},t_{2}\right)$ if and only if $\exists 1\leq\hat{r}\leq n$ is even, s.t.
$\begin{dcases}{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0,\\ \forall 1\leq r<\hat{r},\,{\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)=0,\\ {\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)<0.\end{dcases}$ (29)
@@
Assume that $\forall 1\leq r\leq n$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$. By Lemma 2, $\forall t\in B_{\varepsilon}\left(t_{1}\right)$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)+d\equiv 0$, which contradicts the tangent condition. Hence, $\exists 1\leq\hat{r}\leq n$, s.t. $\forall 1\leq r<\hat{r}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$, and $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}\not=0$.
@@
No matter the arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}\right)$ is constrained or unconstrained, ${\mathcal{S}}$ can be tangent to the boundary of the state constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$. It is noteworthy that ${\mathcal{S}}$ can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$. For this case, ${\mathcal{S}}$ should be constrained. Arbitrarily consider $p\in\mathcal{P}$. According to Proposition 2, ±$\pm u\leq{u_{\mathrm{m}}}$ is equivalent to $\mp\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\leq{u_{\mathrm{m}}}$. In other words, the constraints on $u$ are equivalent to the constraints on ${\boldsymbol{x}}$. Hence, Proposition 5 can apply for both state constraints and control constraints. For convenience of discussion, $\forall 1\leq p\leq P$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$. The $\left(P+1\right)$-th constraint refers to $u\leq{u_{\mathrm{m}}}$, and the $\left(P+2\right)$-th constraint refers to $-u\leq{u_{\mathrm{m}}}$.
@@
Assume that $\forall p\in\mathcal{P}^{\prime}\subset\left[P+2\right]$, an arc ${\mathcal{S}}$ is tangent to the $p$-th constraint at $t_{1}$. Denote the tangent marker as ${\mathcal{T}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}^{\prime}\right)$. $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$ are induced by (29), where $\widehat{{\boldsymbol{C}}}$ has full row rank.
@@ IV-C Augmented Switching Law
Assume that the optimal trajectory of problem (1a) satisfies Assumption 1; hence, it contains a finite number of arcs. The augmented switching law is defined as ${\mathcal{L}}={\mathcal{E}}_{0}\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{k}^{\left(1\right)}\right\}_{k=1}^{M_{1}}\right){\mathcal{E}}_{1}\dots\mathcal{E}_{N-1}\left({\mathcal{S}}_{N},\left\{{\mathcal{T}}_{k}^{\left(N\right)}\right\}_{k=1}^{M_{N}}\right){\mathcal{E}}_{N}$. $\forall i\in\left[N\right]$, ${\mathcal{S}}_{i}$ occurs on $\left[t_{0}^{\left(i\right)},t_{0}^{\left(i+1\right)}\right]$. $\forall k\in\left[M_{i}\right]$, the tangent marker ${\mathcal{T}}_{k}^{\left(i\right)}$ occurs at $t_{k}^{\left(i\right)}$ during ${\mathcal{S}}_{i}$, where $t_{0}^{\left(i\right)}<t_{1}^{\left(i\right)}<t_{2}^{\left(i\right)}\dots<t_{M_{i}}^{\left(i\right)}<t_{0}^{\left(i+1\right)}$. The additional end-constraint ${\mathcal{E}}_{i}$ is induced by Proposition 4 at $t_{0}^{\left(i\right)}$. Among them, ${\mathcal{E}}_{N}$ additionally contains the constraints ${\boldsymbol{x}}\left(t_{0}^{\left(N+1\right)}\right)={{\boldsymbol{x}}_{\mathrm{f}}}$, where $t_{0}^{\left(N+1\right)}={t_{\mathrm{f}}}$. $\forall i\in\left[N+1\right]$, $k\in\left[M_{i}\right]\cup\left\{0\right\}$, the points $t_{k}^{\left(i\right)}$ are called keypoints for the feasibility of the trajectory.
@@
If a feasible trajectory with a finite number of arcs follows the Bang-Singular-Bang control law (7), then Proposition 3, Proposition 4, and Proposition 5 hold for the trajectory. For this case, the augmented switching law can also represent the feasible trajectory.
@@
If a non-chattering Bang-Singular-Bang feasible trajectory is disturbed by a sufficiently small perturbation, then the disturbed trajectory can only exceed the constrained boundary at the keypoints since the distance between the original trajectory and the constrained boundary is 0 at the keypoints, while it is strictly positive at other points. Therefore, the augmented switching law can not only characterize the input control but also describe the feasibility of the trajectory.
@@
The switching law in Definition 2 corresponding to the augmented switching law in Definition 5 is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$. It can be observed that the augmented switching law provides more detailed information on the feasibility of the trajectory at keypoints than the switching law. Notably, neither the switching law nor the augmented switching law provides the motion time of each arc.
To simplify the notation, denote an arc without tangent markers $\left({\mathcal{S}},\varnothing\right)$ as ${\mathcal{S}}$ in an augmented switching law. If no additional end-constraint is induced at the end of an arc, i.e., ${\mathcal{E}}=\left(\sim,\sim,\sim,\sim\right)$, then ${\mathcal{E}}$ is omitted.
For convenience of understanding, the chain-of-integrators system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [8, 18]. The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively. The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively. Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall 1\leq j<k$, $x_{j}\equiv 0$. The tangent marker where ${\boldsymbol{x}}$ is tangent to $x_{{\mathrm{m}}k}$ and $-x_{{\mathrm{m}}k}$ of an even order $2l<k$ in Proposition 5 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively. The additional end-constraint ${\mathcal{E}}=\left(\left(s_{k},r_{k}\right)\right)_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 4, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $-x_{j}\leq x_{{\mathrm{m}}j}$, respectively. Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc. Based on the above simplified notations, an example is provided to illustrate the augmented switching law.
@@
Consider a position-to-position problem for 4th order chain-of-integrators systems with full box constraints shown in Fig. 1. A feasible suboptimal trajectory with a Bang-Singular-Bang form, planned by the manifold-intercepted method (MIM) [8], is shown in Fig. 1(a). It can be observed that the augmented switching law is ${\mathcal{L}}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{03}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$.
Figure 1: Two examples of augmented switching laws. For a 4th order chain-of-integrators system with full box state constraints, let ${u_{\mathrm{m}}}=1$, $x_{{\mathrm{m}}1}=1$, $x_{{\mathrm{m}}2}=1.5$, $x_{{\mathrm{m}}3}=4$, $x_{{\mathrm{m}}4}=15$, ${\boldsymbol{x}}_{0}=-x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$, and ${{\boldsymbol{x}}_{\mathrm{f}}}=x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$. (a) A non-chattering feasible trajectory planned by our previous work [8]. (b) A chattering optimal trajectory planned by our previous work [18]. In (b2), $\hat{x}_{3}\left(t\right)=\left(x_{{\mathrm{m}}3}-x_{3}\left(t\right)\right)\left(t_{\infty}-t\right)^{-3}$ where $t_{\infty}\approx 6.0732$ is a chattering limit time.
@@
Our previous work [18] pointed out that the chattering phenomenon occurs in the optimal trajectory for the problem in Example 3. The chattering optimal trajectory is shown in Fig. 1(b1). Fig. 1(b2) plots the trajectory of $x_{3}$ during the chattering period with amplitude compensation applied. It can be observed that the trajectory during the left chattering period consists of a cycle of $\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}$. According to Bellman’s principle of optimality [20], the trajectory in $\left[0,t_{\infty}-\varepsilon\right]$ is also time-optimal, where $\varepsilon>0$ is small enough. The augmented switching of the above sub-arc begins with $\overline{01}\underline{0}\overline{2}\underline{01}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\dots$
V A State-Centric Necessary Condition for Optimal Control of Problem (1a)
The theoretical framework for the augmented switching law of time-optimal control for linear systems is established in Section IV. Based on the established framework, this section provides a state-centric necessary condition for the optimal control of problem (1a). Section V-A discusses the uniqueness of the time-optimal control. Then, given a feasible Bang-Singular-Bang trajectory, Section V-B analyzes the feasibility of the disturbed Bang-Singular-Bang trajectory. Based on the conclusions in Section V-A and Section V-B, Section V-C provides a necessary condition for the optimal trajectory represented by the augmented switching law, which constrains the difference between the switching times and the sum of orders of keypoints in a way.

### v2

@@ IV-B Tangent Markers in An Arc
Proposition 4 and Proposition 5 analyzed the feasibility of an arc near the end points. An arc might be tangent to some constrained boundary, where small perturbations may affect the feasibility of the arc. The following proposition provides a sufficient and necessary condition for the feasibility of the optimal trajectory tangent to constraints’ boundaries.
@@
Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $\left[t_{0},t_{2}\right]$. ${\mathcal{S}}$ is tangent to the boundary of the constraint ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\leq 0$ at $t_{1}\in\left(t_{0},t_{2}\right)$ if and only if $\exists\hat{r}\in\left[n\right]$ is even, s.t.
$\begin{dcases}{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0,\\ \forall r\in\left[\hat{r}-1\right],\,{\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)=0,\\ {\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{\hat{r}-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{b}}}\right)<0.\end{dcases}$ (29)
@@
Assume that $\forall r\in\left[n\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$. By Lemma 1, $\forall t\in B_{\varepsilon}\left(t_{1}\right)$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)+d\equiv 0$, which contradicts the tangent condition. Hence, $\exists\hat{r}\in\left[n\right]$, s.t. $\forall r\in\left[\hat{r}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}=0$, and $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}\not=0$.
@@
No matter an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is constrained or not, ${\mathcal{S}}$ can be tangent to the boundary of the state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$. It is noteworthy that ${\mathcal{S}}$ can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$. For this case, ${\mathcal{S}}$ should be constrained. Arbitrarily consider $p\in\mathcal{P}$. According to Proposition 2, ±$\pm u\leq{u_{\mathrm{m}}}$ is equivalent to $\mp\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\leq{u_{\mathrm{m}}}$. In other words, the constraints on $u$ are equivalent to the constraints on ${\boldsymbol{x}}$. Hence, Proposition 6 can apply for both state constraints and control constraints. For convenience, $\forall p\in\left[P\right]$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$. The $\left(P+1\right)$-th and $\left(P+2\right)$-th one refer to $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$, respectively.
@@
Consider an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ for $t\in\left[t_{0},t_{2}\right]$. Assume that at $t_{1}\in\left(t_{0},t_{2}\right)$, the arc ${\mathcal{S}}$ is tangent to some constraints’ boundaries whose indices constitute a nonempty set $\mathcal{P}^{\prime}\subset\left[P+2\right]$. By Proposition 6, the tangent condition (29) is equivalent to $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank. Then, the tangent marker is defined as ${\mathcal{T}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}},\mathcal{P}^{\prime}\right)$.
@@ IV-C Augmented Switching Law
Consider a feasible BBS trajectory of problem (1a) with $N$ arcs, denoted as ${\mathcal{S}}_{i}$, $i\in\left[N\right]$. During ${\mathcal{S}}_{i}$, $M_{i}$ tangent markers occur, denoted as ${\mathcal{T}}_{k}^{(i)}$, $k\in\left[M_{i}\right]$. At the connection of ${\mathcal{S}}_{i}$ and ${\mathcal{S}}_{i+1}$, the additional end-constraint is denoted as ${\mathcal{E}}_{i}$. Then, time points at tangent markers and ends of arcs are called keypoints for the trajectory’s feasibility. The augmented switching law (ASL) ${\mathcal{L}}$ is defined in (28).
@@
If a feasible trajectory with a finite number of arcs follows the BBS control law (7), then Propositions 4, 5, and 6 hold for the trajectory. For this case, the ASL can also represent the feasible trajectory.
@@
If a feasible non-chattering BBS trajectory is perturbed by a sufficiently small perturbation, then the perturbed trajectory can only exceed constraints’ boundaries at the keypoints since the distance between the original trajectory and constraints’ boundaries is 0 at the keypoints, while it is strictly positive at other points. Therefore, the ASL can not only characterize the input control but also describe the feasibility of the trajectory.
@@
The switching law in Definition 2 corresponding to the ASL in Definition 5 is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$. It can be observed that the ASL provides more detailed information on the feasibility of the trajectory at keypoints than the switching law. Notably, neither the switching law nor the ASL provides the motion time of each arc.
To simplify the notation, denote an arc without tangent markers $\left({\mathcal{S}},\varnothing\right)$ as ${\mathcal{S}}$ in an ASL. If ${\mathcal{E}}$ is empty, then ${\mathcal{E}}$ is omitted. The additional end-constraints before ${\mathcal{S}}_{1}$ and after ${\mathcal{S}}_{N}$ are omitted since the initial and terminal states are given, i.e., ${\boldsymbol{x}}_{0}$ and ${\boldsymbol{x}}_{\mathrm{f}}$. For the example shown in Fig. 1, the ASL is $\left({\mathcal{S}}_{1},\left\{{\mathcal{T}}_{1}^{(1)}\right\}\right){\mathcal{E}}_{1}\left({\mathcal{S}}_{2},\left\{{\mathcal{T}}_{2}^{(1)},{\mathcal{T}}_{2}^{(2)}\right\}\right){\mathcal{S}}_{3}$.
For convenience of understanding, the chain-of-integrator system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [20, 32]. The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively. The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively. Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$. The tangent marker where ${\boldsymbol{x}}$ is tangent to $x_{{\mathrm{m}}k}$ and $-x_{{\mathrm{m}}k}$ of even order $2l<k$ in Proposition 6 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively. The additional end-constraint =${\mathcal{E}}=\left\{\left(s_{k},r_{k}\right)\right\}_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 5, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $-x_{j}\leq x_{{\mathrm{m}}j}$, respectively. Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc. Based on the above simplified notations, an example is provided to illustrate the ASL.
@@
Consider a position-to-position problem for 4th-order chain-of-integrator systems with full box constraints shown in Fig. 2. A feasible suboptimal trajectory with a BBS form, planned by the manifold-intercepted method (MIM) [20], is shown in Fig. 2(a). It can be observed that the ASL is ${\mathcal{L}}=\overline{01}\underline{0}\overline{2}\underline{01}\overline{03}\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}$.
Figure 2: Two examples of ASLs. For a 4th-order chain-of-integrator system with full box state constraints, let ${u_{\mathrm{m}}}=1$, $x_{{\mathrm{m}}1}=1$, $x_{{\mathrm{m}}2}=1.5$, $x_{{\mathrm{m}}3}=4$, $x_{{\mathrm{m}}4}=15$, ${\boldsymbol{x}}_{0}=-x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$, and ${{\boldsymbol{x}}_{\mathrm{f}}}=x_{{\mathrm{m}}4}{\boldsymbol{e}}_{4}$. (a) A non-chattering feasible trajectory planned by [20]. (b) A chattering optimal trajectory planned by [32]. (b2) is the plot of $\hat{x}_{3}\left(t\right)=\left(x_{{\mathrm{m}}3}-x_{3}\left(t\right)\right)\left(t_{\infty}-t\right)^{-3}$ where $t_{\infty}\approx 6.0732$ is a chattering limit time.
@@
Our previous work [32] pointed out that the chattering phenomenon occurs in the optimal trajectory for the problem in Example 3. The chattering optimal trajectory is shown in Fig. 2(b1). Fig. 2(b2) plots the trajectory of $x_{3}$ during the chattering period with amplitude compensation applied. It can be observed that the trajectory during the left chattering period consists of a cycle of $\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}$. According to Bellman’s principle of optimality [22], the trajectory in $\left[0,t_{\infty}-\varepsilon\right]$ is also time-optimal, where $\varepsilon>0$ is sufficiently small. The augmented switching of the above sub-arc begins with $\overline{01}\underline{0}\overline{2}\underline{01}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\left(\overline{0},\left\{\left(\overline{3},2\right)\right\}\right)\underline{0}\dots$
V A State-Centric Necessary Condition
The theoretical framework for the ASL of time-optimal control for linear systems is established in Section IV. Based on the established framework, this section provides a state-centric necessary condition for the optimal control of problem (1a). Section V-A discusses the uniqueness of the time-optimal control. Then, given a feasible BBS trajectory, Section V-B analyzes the feasibility of the perturbed BBS trajectory. Based on the conclusions in Sections V-A and V-B, Section V-C provides a necessary condition of optimality.